Car goes 30mph in 45 minutes, how far will go in 15?
Found 2 solutions by jim_thompson5910, Edwin McCravy:
Answer by jim_thompson5910(35256)

If you meant 15 minutes, then

d = unknown
r = 45 mph
t = 15 min = (15/60) hrs = 0.25 hrs


d = rt
d = 45*0.25
d = 11.25

The car will go 11.25 miles in 15 minutes

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------

Answer by Edwin McCravy(20081)

15 minutes is one third of 45 minutes. So one-third of 30 miles is 10 miles.